OpenStudy (anonymous):

f(x) = 3x + 6, g(x) = 2x^2 Find (fg)(x). please help

OpenStudy (tkhunny):

What is f(3)?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (tkhunny):

Evaluate f(3).

OpenStudy (anonymous):

so 3=3x+6 evaluate that ?

OpenStudy (tkhunny):

f(3) = 3(3) + 6 -- Perfect. Harder one. Evaluate f(Fred)

OpenStudy (anonymous):

how do i do that one ?

OpenStudy (tkhunny):

Exactly the same way. This was my point. f(Fred) = 3(Fred) + 6 Do you see that this is so?

OpenStudy (anonymous):

u just join it is 6x^3+12x^2

OpenStudy (tkhunny):

Don't get ahead of me. First, though, what does (fg)(x) mean? Is it f(g(x)) or f(x)*g(x)?

OpenStudy (tkhunny):

Anyway, f(g(x)) = 3(g(x)) + 6 = 3(2x^2) + 6 = 6x^2 + 6